# Sparse Hanson-Wright Inequalities with Applications

### v3

@@ 3.1 Sparse Wigner matrices: local law and eigenvector delocalization
Denote the eigenvalues of $H$ as $\lambda_{1}(H),\cdots,\lambda_{n}(H)$ and their corresponding $\ell_{2}$-normalized eigenvectors as $u_{1},\cdots,u_{n}$. When $np\to\infty$, it is known that the empirical spectral distribution (ESD) of $H$, defined as $\mu_{n}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{\lambda_{i}(H)}$, converges weakly to the semicirle law in probability [62, 58]. When $\xi_{ij}=1$, $H$ is the adjacency matrix of Erdős-Rényi graph, and its spectral properties, including the local law and eigenvector delocalization [58, 34], universality [27, 28, 35], and extreme eigenvalues [12, 5], have received considerable attention in recent years. When $\xi_{ij}$ is a bounded random variable, several extensions have been obtained in [5, 57, 8, 26]. When $p=\frac{c}{n}$, the large deviation principle of the spectral edge was studied in [29, 30].
We establish the local law and complete eigenvector delocalization for $H$, analogous to He, Knowles, and Marcozzi’s work [34] that studied the (normalized) adjacency matrix of the Erdős-Rényi random graph $G(n,p)$ in the supercritical regime $np\geq C\log n$. As shown in [6], this is the optimal sparsity level (up to a constant factor) for complete eigenvector delocalization. To the best of our knowledge, this is the first local law and complete delocalization result for sparse $\alpha$-subexponeitial random matrices down to the near-optimal sparsity $p\geq\frac{\mathrm{polylog}(n)}{n}$ and for unbounded sparse sub-gaussian random matrices down to the optimal sparsity $p\geq\frac{C\log n}{n}.$
@@
Another consequence of the local law is the complete delocalization of the (normalized) eigenvectors $\{u_{i}\}_{i=1}^{n}$ of $H$. Note that compared to Theorem  6, we require a slightly stronger lower bound on $p$ due to a technical condition on the spectral norm concentration of $H$ (see Lemma B.1).
THEOREM 8 (Complete eigenvector delocalization for sparse Wigner matrix with $\alpha$-subexponential entries).
@@
THEOREM 9 (Complete eigenvector delocalization for sparse sub-Gaussian random matrices).
@@ 3.2 Concentration of the Euclidean norm of sparse random vectors
Consider $X=(X_{1},\dots,X_{n})^{T}$ a random vector with independent centered sparse $\alpha$-subexponential components, where each $X_{i}$ satisfies $\mathbb{E}(|X_{i}|^{r})\leq p_{i}(Kr^{\frac{1}{\alpha}})^{r}$ for all $r\geq 1$ and some $\alpha\in(0,1]$. Let $B\in\mathbb{R}^{m\times n}$ be an deterministic matrix with columns $B_{1},\dots,B_{n}$. For any $s>0$, with probability at least $1-C^{\prime}\exp(-cs)$, the following holds:
@@
Let us compare our result with the existing bound in [32] for the special case $\alpha=1$, although both their result and ours apply to $\alpha\in(0,1]$. Consider a sparse subexponential random vector $X$ with i.i.d. entries $X_{i}=x_{i}\xi_{i}$ where all $x_{i},\xi_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p)$ and $\xi_{i}$ is a mean-zero, unit-variance random variable with $\|\xi_{i}\|_{L_{r}}\leq Kr$ for all $r\geq 1$. These conditions imply $\|X_{i}\|_{\Psi_{1}}\leq CK$ for a universal constant $C>0$. According to [32, Corollary 2.2], with probability at least $1-C^{\prime}\exp(-cs)$,
@@ 5 Proofs of Theorem 1 and Theorem 2
  • Let $\mathcal{P}(2r)$ be the set of partitions of $[2r]$, and let $\mathcal{P}_{\geq 2}(2r)$ be the set of partitions of $[2r]$ whose blocks are of size at least $2$. For each $\Pi\in\mathcal{P}_{\geq 2}(2r)$, let $s\in[n]^{|\Pi|}$ assign indices $[n]$ to the blocks in $\Pi$. Denote $\mathcal{P}_{\geq 2}(2r,k)\subset\mathcal{P}_{\geq 2}(2r)$ to be the set of partitions of $[2r]$ with exactly $k$ blocks, each of size at least 2. Hence, we focus on $2\leq k\leq r$ since $A$ is diagonal-free and $k=1$ gives a zero in the sum.
  • Encode each $\Pi\in\mathcal{P}_{\geq 2}(2r,k)$ as a multigraph $G(\Pi)$ whose vertex set is $\Pi$. More precisely, consider $\mathcal{G}=([2r],E)$ to be the graph with vertex set $[2r]$ and edge set
@@
    A multigraph $G(\Pi)$ is obtained by identifying the vertices in $\mathcal{G}$ according to the partition $\Pi$. The multigraph $G(\Pi)$ has $r$ edges and $k$ vertices, and each vertex has a degree of at least 2. Also, we focus on the ones that are loop-free, since $A$ is diagonal-free and any loop in $G(\Pi)$ gives a zero entry in the sum.
  • We further denote $\mathcal{P}_{\geq 2}(2r,k,t)\subset\mathcal{P}_{\geq 2}(2r,k)$ to be the set of partitions of $[2r]$ with exactly $k$ blocks and the corresponding graph $G(\Pi)$ has $t$ connected components. Here we only consider $21\leq t\leq k/$ 2, since we assume $A$ is diagonal-free and each connected component contains at least two vertices.
@@
$(eK)^{2r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}\sum_{\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)}\left(\prod_{\tau\in\Pi}\deg_{G(\Pi)}(\tau)!\sum_{s\in[n]^{\Pi}}\left(\prod_{\{\pi_{1},\pi_{2}\}\in E(G(\Pi))}a_{s_{\pi_{1}}s_{\pi_{2}}}\prod_{\pi\in\Pi}p_{s_{\pi}}\right)\right).$ (5.3)
@@
The proof of Lemma 5.2 is deferred to Section 5.1.1. It is worth noting that the bound on $\mathcal{S}_{G(\Pi)}$ in Lemma 5.2 is uniform and independent of $G(\Pi)$. To proceed from (5.5), we now need to estimate
$\sum_{\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)}\prod_{\tau\in\Pi}\deg_{G(\Pi)}(\tau)!.$
@@
    First we consider all the partitions that partition $[2r]$ into $k$ labeled subsets, i.e. each partition is indeed a surjective map $\phi:[2r]\to[k]$. Let $\Phi(\Pi)$ denote the set of such surjection $\phi$ that corresponds to partition $\Pi$. Therefore, as the parts in partition $\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)$ is not ordered, $|\Phi(\Pi)|=k!$. Hence,
@@
    For fixed series $(r_{1},\dots,r_{t})$ and $(k_{1},\dots,k_{t})$, consider the connected components in $G$ are ordered in lexicographic order on the edge partition and the $i$-th component has $r_{i}$ edges and $k_{i}$ vertices. As the order is determined by the edge partition, the edge with the smallest index in the $i$-th component is determined after fixing the edges in the first $i-1$ components. Therefore, there are at most
@@
    Finally, inside $i$-th connected component, there are $r_{i}$ edges and $k_{i}$ labeled vertices $\{j_{1},\dots,j_{k_{i}}\}$. Let us further fix the degree $d_{1},\dots,d_{k_{i}}$ of each vertex and then sum over all possible degree sequences. As the vertices are labeled and the degree of each vertex is fixed, there are at most $\binom{2r_{i}}{d_{1},\dots,d_{k_{i}}}$ many different partition $\phi_{i}$ of $[2r]$. Therefore, with $d_{1}+\dots+d_{k_{i}}=2r_{i}$, we have
@@
Here, the last inequality follows from Lemma A.2 and that the maximum of $\prod_{i}z_{i}^{z_{i}}$ is attained when $t-1$ of the $z_{i}$’s are equal to 1, and the remaining one equals $Z-(t-1)$. Therefore, we have
@@
The main idea for the estimates is to delete the edges in $G=G(\Pi)$ as the following steps:
@@
    For each connected component $\mathcal{C}$ of $G$, we identify a spanning tree $T$ and add one arbitrarily chosen edge from $E(G)\setminus E(T)$, creating a cycle. We then delete all other edges in $\mathcal{C}$. Let $G^{\prime}$ be the remaining subgraph after deletion to all connected components, and hence $|E(G^{\prime})|=k$.
    For each deleted edge, say $\{u,v\}$, we bound the corresponding $a_{uv}\leq\|A\|_{\max}$, and we will delete $r-k$ many edges this way in total, which implies that we delete a total of $r-k$ edges from $G$. These deleted edges collectively contribute a factor of $\|A\|_{\max}^{r-k}$ to bound $\mathcal{S}_{G}$, i.e.,
@@
    where $B:=\sqrt{P}A\sqrt{P}$ and $P:=\mathrm{diag}(p_{1},\dots,p_{n})$. Next, we show that $\|B\|\leq\gamma_{1,\infty}$. To see this, assume that $\|B\|$ is attained at some vector $x=(x_{1},\dots,x_{n})^{T}\in\mathbb{S}^{n-1}$, i.e., $\|B\|=\|Bx\|$. By Cauchy-Schwartz inequality, we have
@@
In the process above, we have $r-k$ edges deleted in the first step and $\ell$ edges deleted in the second step. Then
@@
$\mathbb{E}|X^{T}AX|^{r}\leq(eK)^{2r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}\sum_{\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)}\left(\prod_{\tau\in\Pi}\deg_{G(\Pi)}(\tau)!\deg_{G(\Pi)}(\tau)^{(\frac{1}{\alpha}-1)\deg_{G(\Pi)}(\tau)}\mathcal{S}_{G(\Pi)}\right),$ (5.38)
@@
We claim that $\mathcal{D}(\Pi)\leq 2^{2r}r^{2(r-k+1)}$ uniformly for all $\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)$. Consider any fixed graph $G(\Pi)$ with $r$ edges and $k$ vertices. Let its degree sequence be denoted by $d_{1},\dots,d_{k}$. We have $d_{i}\geq 2$ and $d_{1}+\dots+d_{k}=2r$. Therefore, by Lemma A.2
@@ 5.4 Proof of Theorem 2
=$\underset{(\text{c})}{=}K^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\frac{1}{k!}\sum_{\begin{subarray}{c}\phi:[r]\to[k],\\ |\phi^{-1}(i)|\geq 2\end{subarray}}\left(\prod_{i=1}^{k}|\phi^{-1}(i)|^{|\phi^{-1}(i)|}\right)^{\frac{1}{\alpha}}$ (5.55)
@@
6 Proofs of Theorems 4 and 5

### v4

@@ 3.1 Sparse Wigner matrices: local law and eigenvector delocalization
Denote the eigenvalues of $H$ as $\lambda_{1}(H),\cdots,\lambda_{n}(H)$ and their corresponding $\ell_{2}$-normalized eigenvectors as $u_{1},\cdots,u_{n}$. When $np\to\infty$, it is known that the empirical spectral distribution (ESD) of $H$, defined as $\mu_{n}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{\lambda_{i}(H)}$, converges weakly to the semicircle law in probability [68, 64]. When $\xi_{ij}=1$, $H$ is the adjacency matrix of Erdős-Rényi graph, and its spectral properties, including the local law and eigenvector delocalization [64, 37], universality [29, 30, 39], and extreme eigenvalues [13, 6], have received considerable attention in recent years. When $\xi_{ij}$ is a bounded random variable, several extensions have been obtained in [6, 63, 10, 28]. When $p=\frac{c}{n}$, the large deviation principle of the spectral edge was studied in [31, 32].
We establish the local law and complete eigenvector delocalization for $H$, analogous to He, Knowles, and Marcozzi’s work [37] that studied the (normalized) adjacency matrix of the Erdős-Rényi random graph $G(n,p)$ in the supercritical regime $np\geq C\log n$. As shown in [7], this is the optimal sparsity level (up to a constant factor) for complete eigenvector delocalization. To the best of our knowledge, this is the first local law and complete delocalization result for sparse $\alpha$-subexponential random matrices down to the near-optimal sparsity $p\geq\frac{\mathrm{polylog}(n)}{n}$ and for unbounded sparse sub-gaussian random matrices down to the optimal sparsity $p\geq\frac{C\log n}{n}.$
@@
Another consequence of the local law is the complete delocalization of the (normalized) eigenvectors $\{u_{i}\}_{i=1}^{n}$ of $H$. Note that compared to Theorem  4, we require a slightly stronger lower bound on $p$ due to a technical condition on the spectral norm concentration of $H$ (see Lemma B.1).
THEOREM 6 (Eigenvector delocalization for sparse Wigner matrix with $\alpha$-subexponential entries).
@@
THEOREM 7 (Eigenvector delocalization for sparse sub-Gaussian random matrices).
@@ 3.2 Concentration of the Euclidean norm of sparse random vectors
Consider $X=(X_{1},\dots,X_{n})^{T}$ a random vector with independent centered sparse $\alpha$-subexponential components, where each $X_{i}$ satisfies $\mathbb{E}(|X_{i}|^{r})\leq p_{i}(Kr^{\frac{1}{\alpha}})^{r}$ for all $r\geq 1$ and some $\alpha\in(0,1]$. Let $B\in\mathbb{R}^{m\times n}$ be a deterministic matrix with columns $B_{1},\dots,B_{n}$. For any $s>0$, with probability at least $1-C^{\prime}\exp(-cs)$, the following holds:
@@
Let us compare our result with the existing bound in [35] for the special case $\alpha=1$, although both their result and ours apply to $\alpha\in(0,1]$. Consider a sparse subexponential random vector $X$ with i.i.d. entries $X_{i}=x_{i}\xi_{i}$ where all $x_{i},\xi_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p)$ and $\xi_{i}$ is a mean-zero, unit-variance random variable with $\|\xi_{i}\|_{L_{r}}\leq Kr$ for all $r\geq 1$. These conditions imply $\|X_{i}\|_{\Psi_{1}}\leq CK$ for a universal constant $C>0$. According to [35, Corollary 2.2], with probability at least $1-C^{\prime}\exp(-cs)$,
@@ 5 Proofs of Theorem 1 and Theorem 2
  • Let $\mathcal{P}(2r)$ be the set of partitions of $[2r]$, and let $\mathcal{P}_{\geq 2}(2r)$ be the set of partitions of $[2r]$ whose blocks are of size at least $2$. For each $\Pi\in\mathcal{P}_{\geq 2}(2r)$, let $s\in[n]^{|\Pi|}$ assign indices $[n]$ to the blocks in $\Pi$. Denote $\mathcal{P}_{\geq 2}(2r,k)\subset\mathcal{P}_{\geq 2}(2r)$ to be the set of partitions of $[2r]$ with exactly $k$ blocks, each of size at least 2.
  • Encode each $\Pi\in\mathcal{P}_{\geq 2}(2r,k)$ as a multigraph $G(\Pi)$ whose vertex set is $\Pi$. More precisely, consider $\mathcal{G}=([2r],E(\mathcal{G}))$ to be the graph with vertex set $[2r]$ and edge set
@@
    A multigraph $G(\Pi)$ is obtained by defining $V(G(\Pi))=\Pi$ and
    $E(G(\Pi))=\{\{\tau_{2\ell-1},\tau_{2\ell}\}\mid\ell\in[r];\;i\in\tau_{i}\in\Pi\;\forall i\in[2r]\}.$
    The multigraph $G(\Pi)$ has $r$ edges and $k$ vertices, and each vertex has a degree of at least 2. For each vertex $\tau\in\Pi$, let $\deg_{G(\Pi)}(\tau)$ be the degree of $\tau$ in $G(\Pi)$.
  • We further denote $\mathcal{P}_{\geq 2}(2r,k,t)\subset\mathcal{P}_{\geq 2}(2r,k)$ to be the set of partitions of $[2r]$ with exactly $k$ blocks and the corresponding graph $G(\Pi)$ has $t$ connected components.
@@
$(eK)^{2r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}\sum_{\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)}$ $\Bigg(\prod_{\tau\in\Pi}\deg_{G(\Pi)}(\tau)!$
$\qquad\times\sum_{s\in[n]^{\Pi}}\left(\prod_{\{\pi_{1},\pi_{2}\}\in E(G(\Pi))}a_{s_{\pi_{1}}s_{\pi_{2}}}\prod_{\pi\in\Pi}p_{s_{\pi}}\right)\Bigg).$ (5.3)
Here in the sum, the variable $\Pi$ is a partition of $[2r]$, which indicates what group of $X_{i_{1}},\dots,X_{i_{2r}}$ are the copies of the same random variable. And $s:\Pi\to[n]$ indicates the index of $X_{1},\dots,X_{n}$ that a certain group in $\Pi$ corresponds to. Note that $s$ should be an injection: if $i=s(\tau_{1})=s(\tau_{2})$ for some $\tau_{1},\tau_{2}\in\Pi$, by definition $X_{i_{j_{1}}}$ and $X_{i_{j_{2}}}$ are the copy of the same random variable $X_{i}$, $\forall j_{1}\in\tau_{1},j_{2}\in\tau_{2}$, and hence $\tau_{1}=\tau_{2}$ are the same group. We bound (5.2) with (5.1) by changing the sum over $s$ from all possible injections $\Pi\to[n]$ to all maps $n]^{\Pi}$.
Also, note that $A$ is diagonal-free. If there is some loop edge $\{\pi_{1},\pi_{2}\}\in E(G(\Pi))$ with ${\pi_{1}}={\pi_{2}}$, there will be $a_{s_{\pi_{1}}s_{\pi_{2}}}=0$, annihilating the product and also the whole sum. Therefore, in the later discussion, we will focus on the subset of $\{\Pi\in\mathcal{P}_{\geq 2}(2r,k)\mid G(\Pi)\text{ is loop-free}\}$. Note that it also implies $2\leq k\leq r$ as it is impossible to be loop-free when $k=1$.
@@
The proof of Lemma 5.2 is deferred to Section 5.1.1. It is worth noting that the bound on $\mathcal{S}_{G(\Pi)}$ in Lemma 5.2 is uniform and independent of $G(\Pi)$. To proceed from (5.5), we claim that the following bound holds:
$\sum_{\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)}\prod_{\tau\in\Pi}\deg_{G(\Pi)}(\tau)!\leq 2^{6r}e^{(3+2/e)r}r^{2r-k-t}.$ (5.7)
We establish (5.7) in Step 3 below.
@@
    First we consider all the partitions that partition $[2r]$ into $k$ labeled subsets, i.e. each partition is indeed a surjective map $\phi:[2r]\to[k]$. Let $\Phi(\Pi)$ denote the set of such surjection $\phi$ that corresponds to partition $\Pi$. Therefore, as the parts in partition $\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)$ are not ordered, $|\Phi(\Pi)|=k!$. Hence,
@@
    For fixed series $(r_{1},\dots,r_{t})$ and $(k_{1},\dots,k_{t})$, consider the connected components in $G$ are ordered in lexicographic order on the edge partition and the $i$-th component has $r_{i}$ edges and $k_{i}$ vertices. As the order is determined by the edge partition, the edge with the smallest index in the $i$-th component is determined after fixing the edges in the first i1i-1 components. Therefore, there are at most
@@
    Finally, inside $i$-th connected component, there are $r_{i}$ edges and $k_{i}$ labeled vertices $\{j_{1},\dots,j_{k_{i}}\}$. Let us further fix the degree $d_{1},\dots,d_{k_{i}}$ of each vertex and then sum over all possible degree sequences. As the vertices are labeled and the degree of each vertex is fixed, there are at most $\binom{2r_{i}}{d_{1},\dots,d_{k_{i}}}$ many different partitions $\phi_{i}$ of $2r_{i}]$. Therefore, with $d_{1}+\dots+d_{k_{i}}=2r_{i}$, we have
@@
Here, the last inequality follows from Lemma A.2 and that the maximum of $\prod_{i}z_{i}^{z_{i}}$ is attained when t1t-1 of the $z_{i}$’s are equal to 1, and the remaining one equals Z(t1)Z-(t-1). Therefore, we have
@@
The main idea for the estimates is to delete the edges in $G=G(\Pi)$ as follows:
@@
    For each connected component $\mathcal{C}$ of $G$, we consider any fixed spanning subgraph $H_{\mathcal{C}}$ in $\mathcal{C}$ that contains exactly one cycle (or equivalently, spanning subgraph with $|E(H_{\mathcal{C}})|=|V(H_{\mathcal{C}})|$; such structure can be obtained by first picking any fixed spanning tree of $\mathcal{C}$ and then adding one extra edge to it). We then delete all other edges in $\mathcal{C}$. Let $G^{\prime}$ be the remaining subgraph (the union of all $H_{\mathcal{C}}$) after the deletion of all connected components, and hence $|E(G^{\prime})|=|V(G^{\prime})|=k$.
    For each deleted edge, say $\{u,v\}$, we bound the corresponding $a_{uv}\leq\|A\|_{\max}$, and we will delete rkr-k many edges this way in total, which implies that we delete a total of rkr-k edges from $G$. These deleted edges collectively contribute a factor of $\|A\|_{\max}^{r-k}$ to bound $\mathcal{S}_{G}$, i.e.,
@@
    where $B:=\sqrt{P}A\sqrt{P}$ and $P:=\mathrm{diag}(p_{1},\dots,p_{n})$. Next, we show that $\|B\|\leq\gamma_{1,\infty}$. To see this, assume that $\|B\|$ is attained at some vector $x=(x_{1},\dots,x_{n})^{T}\in\mathbb{S}^{n-1}$, i.e., $\|B\|=\|Bx\|$. By the Cauchy-Schwarz inequality, we have
@@
In the process above, we have rkr-k edges deleted in the first step and $\ell$ edges deleted in the second step. Then
@@
$\mathbb{E}|X^{T}AX|^{r}$
$\leq(eK)^{2r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}\sum_{\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)}\!\left(\prod_{\tau\in\Pi}\deg_{G(\Pi)}(\tau)!\deg_{G(\Pi)}(\tau)^{(\frac{1}{\alpha}-1)\deg_{G(\Pi)}(\tau)}\mathcal{S}_{G(\Pi)}\!\right),$ (5.39)
@@
We claim that $\mathcal{D}(\Pi)\leq 2^{2r}r^{2(r-k+1)}$ uniformly for all $\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)$ due to Lemma A.2. Consider any fixed graph $G(\Pi)$ with $r$ edges and $k$ vertices. Let its degree sequence be denoted by $d_{1},\dots,d_{k}$. We have $d_{i}\geq 2$ and $d_{1}+\dots+d_{k}=2r$. Therefore, by Lemma A.2,
@@ 5.4 Proof of Theorem 2
$\underset{(\text{c})}{=}K^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\frac{1}{k!}\sum_{\begin{subarray}{c}\phi:[r]\to[k],\\ |\phi^{-1}(i)|\geq 2\end{subarray}}\left(\prod_{i=1}^{k}|\phi^{-1}(i)|^{|\phi^{-1}(i)|}\right)^{\frac{1}{\alpha}}$ (5.56)
@@
6 Improved precise tail bounds for the linear and quadratic forms
Before presenting our improved, precise tail bounds for the linear and quadratic forms for $\alpha$-subexponential random variables with $\alpha\in(1,\infty]$, we introduce several essential notations.
Throughout the context, we denote
$\beta:=1-\frac{1}{\alpha}\quad\text{for }\alpha\in(1,\infty].$ (6.1)
THEOREM 9 (Tail bound for the linear form: $\alpha>1$).
where we denote
REMARK 6.1.
$\exp\left(-C_{\alpha}\frac{t}{KM}\cdot H_{\alpha}(T)\right).$
Next, we present the improved tail bound for the quadratic forms.
THEOREM 10 (Tail bound for the quadratic form: $\alpha>1$).
where we denote
with
REMARK 6.2.
Due to the implicit nature of the Lambert W function, for practical applications of Theorems 9 and 10, we utilize the bounds established in [22]:
$-1-\sqrt{2u}-u<W_{-1}(-e^{-u-1})<-1-\sqrt{2u}-\frac{2}{3}u$ (6.5)
for $u>0$. One can readily verify that Theorem 3 follows directly from applying these bounds (6.5) to Theorems 9 and 10, observing that $W_{-1}\big(-\frac{1}{\beta}T^{-\frac{1}{\beta}}\big)=W_{-1}\big(-e^{-\log\big(\frac{\beta}{e}T^{\frac{1}{\beta}}\big)-1}\big)$.